### v1

# FinSage: A Multi-aspect RAG System for Financial Filings Question Answering

Abstract.
Leveraging large language models in real-world settings often entails a need to utilize domain-specific data and tools in order to follow the complex regulations that need to be followed for acceptable use. Within financial sectors, modern enterprises increasingly rely on Retrieval-Augmented Generation (RAG) systems to address complex compliance requirements in financial document workflows. However, existing solutions struggle to account for the inherent heterogeneity of data (e.g., text, tables, diagrams) and evolving nature of regulatory standards used in financial filings, leading to compromised accuracy in critical information extraction. We propose the FinSage framework as a solution, utilizing a multi-aspect RAG framework tailored for regulatory compliance analysis in multi-modal financial documents. FinSage introduces three innovative components: (1) a multi-modal pre-processing pipeline that unifies diverse data formats and generates chunk-level metadata summaries, (2) a multi-path sparse-dense retrieval system augmented with query expansion (HyDE) and metadata-aware semantic search, and (3) a domain-specialized re-ranking module fine-tuned via Direct Preference Optimization (DPO) to prioritize compliance-critical content. Extensive experiments demonstrate that FinSage achieves an impressive recall of 92.51% on 75 expert-curated questions derived from surpasses the best baseline method on the FinanceBench question answering datasets by 24.06% in accuracy. Moreover, FinSage has been successfully deployed as financial question-answering agent in online meetings, where it has already served more than 1,200 people.
LLM, Information Retrieval, Document Pre-Processing, Large Language Models, Retrieval Augmented Generation, Generative Models, Synthetic Datasets, Generative Retrieval
Figure 1. An overview of our proposed FinSage.
1. Introduction
In the real world, various regulatory constraints define the manners in which tools and systems can be permissibly integrated into modern workflows, complicating the manners in which technologies such as large language models (LLMs) can be directly play a role in fulfilling specific tasks. One such domain is the financial sector, where adherence to regulations is essential to mitigate legal risk and maintain stakeholder confidence. However, financial regulations can evolve with time, presenting complex and significant challenges even for well-structured institutions with expert teams dedicated for this specific purpose (SEC, 2024b, a). More practically, these challenges have led to a strong demand for a Question Answering (QA) system capable of efficiently retrieving and analyzing compliance-related information from financial filings (Shah et al., 2024), with the hope of reducing manual labor and minimizing human error. Such systems often require the integration of retrieval-augmented generation (RAG), which enables machine learning-powered systems to directly search through databases to extract relevant information that can empower greater decision making.
However, while RAG systems have proven helpful for general LLM reliability and reasoning, integration into financial compliance tasks can be difficult due to inherent domain-related limitations. First, financial filings combine unstructured text (narrative disclosures), semi-structured data (tables, graphs), and contextual metadata - a multi-modal composition that conventional text-centric retrieval architectures fail to process cohesively, leading to fragmented or incomplete context representation (Xie et al., 2024; Riedler and Langer, 2024; Zhai, 2024; Xia et al., 2024). Next, while existing RAG systems predominantly employ dense retrieval (Karpukhin et al., 2020; Xiong et al., 2020) or sparse lexical matching (Robertson et al., 2009), these can fail due to the need for domain-specific fine-tuning or an inability to capture implicit regulatory relationships (Jimeno-Yepes et al., 2024; Yuan et al., 2024; Sultania et al., 2024). Finally, existing retrieval methods often rely on ranking relevant passages based on semantic similarity rather than domain-specific criteria, leading to the critical risk of faulty or incomplete reasoning (Barnett et al., 2024; Besta et al., 2024; Agrawal et al., 2024). Such limitations collectively undermine the precision and accountability of automated compliance workflows, exposing enterprises to operational and regulatory hazards.
As a solution to these challenges, we introduce FinSage, an end-to-end financial question-answering framework optimized for regulatory compliance that leverages the heterogeneous nature of financial data for better reasoning. As shown in Figure 1, FinSage consists of three core components: (1) Financial Filings Pre-processing (FFP) pipeline, (2) Multi-path Retrieval (MPR) pipeline, and (3) Domain-Specialized Document Re-ranker, which collectively overcome the aforementioned challenges of multi-modal documents domain-agnostic retrieval, and ranking ambiguity. First, the FFP module tackles the multi-modal heterogeneity in financial documents by unifying text, tables, and graphs into a structured vector database through modality-specific encoders. Next, the MPR module augments traditional dense/sparse retrieval with two tailored domain-aware strategies, ensuring robust coverage of both explicit and implicit domain-specific relationships. Finally, the Document Re-ranker addresses precision limitations through direct preference optimization (DPO) (Zhang et al., 2024a), prioritizing chunks containing legally significant phrases (e.g., “material weakness” or “non-compliance”) and suppressing irrelevant semantic matches. By jointly optimizing retrieval diversity and precision, FinSage establishes a new paradigm for compliance-critical financial QA systems. The experimental results demonstrate that FinSage significantly outperforms traditional RAG approaches in financial regulatory compliance tasks, offering a practical and scalable solution for financial institutions. In summary, our contributions are as follows:
  • We present an end-to-end financial QA system that integrates pre-processing, retrieval, and re-ranking to provide a robust solution for regulatory compliance.
  • We address key challenges in multi-modal financial document processing and domain-specific question answering, bridging the gap between traditional RAG models and regulatory compliance applications.
  • We conduct extensive experiments and ablation studies to evaluate the effectiveness of our retrieval, re-ranking, and response generation mechanisms, analyzing the impact of different retrieval paths and re-ranking strategies.
Figure 2. The complete FinSage RAG pipeline. The system processes the user’s query and splits it into subqueries. For each subquery, multiple chunks are retrieved with multi-path retrieval, and these chunks are then expanded into candidate bundles by concatenating neighboring chunks. After re-ranking, the most relevant chunks and their bundles are included for answer generation.
2. Related Work
2.1. Retrieval Augmented Generation
Despite their strong empirical performance, large language models (LLMs) are inherently bottlenecked by their fixed parametric size, limiting their ability to store factual knowledge, which can affect their performance on specific tasks. Retrieval-Augmented Generation (RAG) (Cai et al., 2022; Ma et al., 2025) emerged as a solution by integrating external knowledge that the LLM could access through retrieval within an external database. A retrieval module first selects relevant documents from sources like news, academic papers, or social media, which are then combined with the input query and fed into a language model to generate a response. This approach leverages both the model’s internal memory and the retrieved corpus (Zhao et al., 2024), ensuring more accurate and contextually grounded outputs.
Recent studies have explored various improvements, such as feedback tokens for relevance (Asai et al., 2024), self-referential knowledge acquisition (Wang et al., 2023), iterative self-feedback (Liu et al., 2024), and domain-specific hierarchies (Wang et al., 2024). In cases where initial queries lack sufficient detail, generative retrieval enhances information retrieval by generating queries or synthesizing passages beyond traditional keyword-based search (Rajput et al., 2023; Li et al., 2024b; Anonymous, 2024). Some studies further work to use multi-head RAG to extract information from diverse sources (Besta et al., 2024; Xia et al., 2024), aiming to maximize answer accuracy through quantity-based retrieval. Others focus on enhancing answer precision via quality retrieval, leveraging re-rankers (Jacob et al., 2024; de Souza P. Moreira et al., 2024; Li et al., 2024a) to preserve high answer relevancy.
2.2. Multi-Modal Retrieval
Information is not limited to text, meaning that knowledge from various modalities can be used for improved reasoning. Some approaches, such as MuRAG (Chen et al., 2022) and generative retrieval frameworks (Long et al., 2024), enhance retrieval by leveraging multi-modal memory, generative retrieval, and knowledge-guided decoding, improving performance in multi-modal QA tasks. Others, including VisRAG (Yu et al., 2025) and multi-modal RAG for industrial applications (Riedler and Langer, 2024), explore vision-language models to retain document structure and improve retrieval accuracy, particularly in industrial and document-heavy settings.
2.3. Retrieval for Financial Data
Despite recent advances, RAG-LLM systems remain underdeveloped for financial filings. Islam et al. (2023) introduced FinanceBench, a dataset specifically designed for financial document understanding. Building on this, Unstructured (Jimeno-Yepes et al., 2024) explores different chunking strategies to enhance financial document processing. Setty et al. (2024) proposes a RAG pipeline to facilitate retrieval in financial reports. However, existing approaches still face significant challenges: they struggle to handle complex multi-modal information—such as figures and tables—and their simple cross-encoder-based re-rankers often lack precision, negatively impacting retrieval performance.
3. Methodology
To reliably use RAG for financial document processing, the system must address three fundamental challenges: (1) handling multi-modal data such as tables, figures, and structured text; (2) extracting a comprehensive scope of relevant information from multiple sources; and (3) generating accurate and precise responses by effectively summarizing and reasoning over retrieved content. To tackle these challenges, we propose FinSage, an end-to-end financial QA system tailored for regulatory compliance.
This section introduces the three key components of FinSage, each designed to address specific challenges in financial RAG (Figure 2). Section 3.1 presents the Financial Filings Pre-processing Pipeline (FFP), which transforms financial documents into a structured machine-readable format while enriching their semantic content. Section 3.2 details the Multi-path Retrieval Pipeline (MPR), which enhances retrieval accuracy and contextual relevance by integrating multiple retrieval strategies. Section 3.3 discusses the Document Re-ranker, which improves retrieval precision by refining and prioritizing the most relevant document chunks.
Figure 3. The FinSage FFP pipeline. Textual Encoding: PDF parsing tools extract multi-modal chunks (text, figures, tables), which are then converted into textual representations using large language models (LLMs). Semantic Enhancement: The textual representations are further refined by (a) eliminating redundant chunks through pairwise similarity comparison, (b) resolving co-reference within subtitle sections, and (c) generating subtitle-section-based summaries as metadata. These enriched chunks are then embedded in a vector database for further processing.
3.1. Financial Filings Pre-processing pipeline
We create a Financial Filings Pre-processing (FFP) pipeline to convert multi-modal financial filings into machine-readable text chunks. Figure 3 illustrates how FFP operates in two steps: textual encoding and semantic enhancement.
Formally, let $\mathcal{C}$ denote the original set of multi-modal chunks extracted from a financial document, divided based on its chapters and a pre-defined token length. Each chunk $s_{i}=\left(t_{i},m_{i}\right)\in\mathcal{C}$ consists of raw content $t_{i}$ and metadata $m_{i}$. The goal of the FFP is to transform chunks in $\mathcal{C}$ into $\tilde{\mathcal{C}}$, a self-contained and condensed format that is better suited for downstream information retrieval tasks.
Textual Encoding of Multi-Modal Documents.
The FFP pipeline first converts multi-modal financial documents into structured, machine-readable textual formats. Following the natural reading order, we use open-source PDF parsing tools, such as MinerU (Contributors, 2024), to decompose each document into a sequence of chunks. Each chunk is labeled with its corresponding type (i.e., text, figure, or table). Inspired by Jimeno-Yepes et al. (2024), we further transform figure and table chunks into descriptive textual narratives with large vision-language models. Specifically, figures are converted into structured captions summarizing their key insights, while tables are rewritten as textual statements capturing their primary data trends and relationships (Zhang et al., 2023). This transformation enables large language models (LLMs) to effectively interpret and utilize multi-modal content.
Semantic Enhancements.
To improve the quality and coherence of the extracted texts, the FFP pipeline applies the following three key semantic enhancement steps.
  1. (1)
    Redundant chunk de-duplication. We first de-duplicate documents to reduce redundancy, ensuring that duplicate content does not limit the diversity of retrieved information. For any two chunks $s_{i},s_{j}\in\mathcal{C}$ such that $i\neq j$, we randomly discard $s_{i}$ or $s_{j}$ if the cosine similarity of the dense embeddings of their raw texts $\cos\left(\mathbf{d}_{i},\mathbf{d}_{j}\right)$ exceeds a predefined threshold $\tau_{\text{dedup}}$.
  2. (2)
    Co-reference resolution. Next, we disambiguate the textual context of each chunk by iteratively resolving co-references. Within each subtitle section, pronouns of entity mentions are replaced with their corresponding antecedents identified from preceding chunks. Specifically, for each chunk containing pronouns, we employ an LLM as a co-reference resolver to perform entity mention replacements. The LLM receives $k$ previous chunks to identify proper antecedents, where $k$ is fixed.
  3. (3)
    Metadata generation. Due to the token-length limitation, document sections may be separated into multiple chunks where content may no longer be self-contained. Thus, we ensure that chunks within a section retain sufficient context from others in the same section for retrieval. We employ an LLM to summarize each section of the financial filing and append the generated summary to the metadata of every chunk obtained from that section. This ensures that each chunk retains enough contextual information while maintaining a minimal context length, enhancing the efficiency and accuracy of document retrieval.
After the aforementioned process, for each chunk $s_{i}$, dense and sparse embeddings $\mathbf{d}_{i}$ and $\mathbf{p}_{i}$ are generated from $\tilde{t}_{i}$ using a text encoder and BM25, respectively. Furthermore, a dense embedding for the metadata $\mathbf{e}_{i}$ is also generated by a text encoder. The chunk is then represented as a 5-tuple $(\tilde{t}_{i},\tilde{m}_{i},\mathbf{d}_{i},\mathbf{p}_{i},\mathbf{e}_{i})$. These enhancements ensure that extracted chunks are self-contained, contextually coherent, and semantically enriched, optimizing their utility for downstream processing. Further details regarding the selection of the multi-modal textualization model and the dense embedding model are found in Appendices A.3 and A.4, respectively.
3.2. Multi-path Retrieval pipeline
We introduce a Multi-path Retrieval (MPR) pipeline that enhances performance by leveraging both metadata and semantic information during the retrieval process. The pipeline consists of two primary modules: a Query Paraphrasing Module and a Retrieval Module for obtaining relevant answers from diverse sources.
Query Paraphrasing Module
An LLM agent is employed to reformulate each input query using FinSage’s specialized vector store, which serves as a comprehensive knowledge base containing all company-related information. This paraphrasing process involves three key steps:
  1. (1)
    Query Decomposition: Each query is segmented into multiple self-contained sub-queries. This ensures that each subquery can be independently addressed by mapping its keywords to the corresponding segments in the knowledge base.
  2. (2)
    Co-reference Resolution: Co-references within each query are resolved by converting the text into standardized English. In this process, all pronouns are replaced with their corresponding antecedents.
  3. (3)
    Context Integration: For queries that reference prior conversation history, relevant contextual information is incorporated to maintain completeness and clarity.
Retrieval Module
Traditional retrieval methods typically rely on either sparse or dense retrieval techniques, each presenting challenges when applied to domain-specific tasks. Dense retrievers generally require extensive fine-tuning with substantial resources, whereas sparse retrievers often depend on heuristic rules that can lead to inaccuracies. To address these limitations, we incorporate both metadata-based and hypothesis-based retrieval strategies, resulting in a multi-path retrieval system. Consequently, our MPR pipeline comprises four components:
  1. (1)
    BM25 Sparse Retriever: This component employs the BM25 scoring function to compute relevance based on tf-idf between the query and documents (Robertson et al., 2009).
  2. (2)
    BGE-M3 Dense Retriever: Utilizing the BGE-M3 model (Chen et al., 2024), this retriever calculates the cosine similarity between query embeddings and chunk embeddings. The retrieval process is accelerated by FAISS (Johnson et al., 2019), ensuring high-performance search.
  3. (3)
    Metadata Retriever: Leveraging the same BGE-M3 model, this module computes the cosine similarity between the query and metadata embeddings, facilitating the integration of structured metadata into the retrieval process.
  4. (4)
    HyDE Retriever: This component computes the cosine similarity between hypothesis and chunk embeddings.
We collect all results selected by those retrievers into a chunk set, which serves as the output of the retrieval module.
Our metadata retrieval approach is designed to address high-level search tasks, such as question answering that requires reasoning across multiple text segments. Traditional retrieval methods often struggle to capture all relevant chunks; therefore, we incorporate metadata summaries (generated via FFP) under a consistent subtitle to enhance retrieval coherence. Specifically, our approach concatenates each chunk’s heading and summary to form a unified metadata embedding. In our framework, every chunk within a semantic segment is assigned an identical heading-summary representation, ensuring that the retrieval of a single metadata instance inherently retrieves all associated chunks. This design not only reinforces the contextual linkage among chunks from the same document segment but also effectively mitigates retrieval ambiguity when dealing with multiple documents.
The HyDE retriever is designed to further address recall issues. In certain scenarios, target chunks and user queries exhibit only weak correlations, even when the queries are optimally formulated. In such cases, the HyDE retriever leverages Hypothetical Document Embeddings(Zhang et al., 2024b), a specialized query expansion technique which uses an LLM-generated response to the user query as a hypothesis for retrieving similar chunks. To further enhance HyDE’s performance, we augment the LLM’s domain-specific knowledge through instruction fine-tuning on our custom Company dataset, following the InPars method (Bonifacio et al., 2022). Additionally, a smaller model is subsequently trained on a synthesized dataset to emulate the generative capabilities of larger models. As a result, HyDE not only improves the matching accuracy between the generated hypothetical responses and the target chunks but also bolsters overall retrieval performance across diverse content types.
Chunk Bundling Module
After retrieval, the MPR pipeline employs a chunk expansion mechanism to enhance contextual coherence among sequential chunks. This approach is particularly critical in financial documents, where key topics and discussions often span multiple chunks or paragraphs. Initially, each chunk is treated as an independent unit. These units are then iteratively expanded by evaluating adjacent chunks (preceding, following, or both) based on a predefined semantic similarity threshold. Formally, for each initial chunk $s_{i}$, we combine it with adjacent chunks $s_{i-1}$ or $s_{i+1}$ if the cosine similarity of their dense embeddings, namely $\cos(\mathbf{d}_{i},\mathbf{d}_{i-1})$ or $\cos(\mathbf{d}_{i},\mathbf{d}_{i-1})$) exceeds an empirically determined threshold $\tau_{\text{exp}}$.
3.3. Document Re-ranking
The Document Re-ranker (DRR) in FinSage acts as an intelligent filter, addressing shortcomings of initial retrieval to enhance accuracy, coherence, and reliability in final generated responses. We employ the BAAI/bge-reranker-v2-Gemma model based on its superior performance in both English and multilingual applications. Within FinSage, the re-ranker refines retrieved documents before passing them to the LLM, providing more relevant contextual input for answer generation. Specifically, the Document Re-ranker scores candidate chunks $s\in\mathcal{R}_{q}$ as $\mathcal{K}\left(s,q\right)$ using a cross-encoder:
(1) $\mathcal{K}\left(s_{i},q\right)=\sigma\left(\mathbf{w}^{\top}\text{CrossEncoder}\left(\text{concat}\left(q,\tilde{t}_{i}\right)\right)+\beta f\left(\tilde{m}_{i}\right)+b\right),$
where $\textbf{w}^{\top}\text{CrossEncoder}(\cdot)+b$ is a Transformer-based re-ranker scoring contextual alignment (Déjean et al., 2024; Askari et al., 2023), and $f\left(\tilde{m}_{i}\right)$ processes metadata to output a scalar bonus, scaled by $\beta$. The final score is transformed using a sigmoid function $\sigma$.
Time bonus
When ranking chunks for relevance, a common challenge arises when two similar chunks lead to favoring older content instead of the new, potentially more desirable entries. Recognizing that metadata, particularly publication dates, is crucial for evaluating chunk relevance, we incorporate a time bonus for chunk $s_{i}$, $f\left(\tilde{m}_{i}\right)$, into the overall ranking score in DRR. Specifically, if the publication date of the chunk is within a year of the query time, there will be an additional bonus. The closer the publication date is to the query time, the larger this bonus will be. This mitigates the issue where the DRR favors older contents over new one when the two chunks have identical embeddings.
Direct Preference Optimization
To further enhance ranking performance, we fine-tune the DRR using Direct Preference Optimization (DPO). This method frames preference learning as a binary classification task, where training data consists of preference pairs: a preferred (positive) response $s^{+}$ and a rejected (negative) response $s^{-}$. The model learns to prioritize high-quality responses by minimizing the cross-entropy loss:
(2) $\mathcal{L}{(\text{DPO})}=-\mathbb{E}{(q,s^{+},s^{-})}\left[\log\frac{\mathcal{K}(s^{+},q)}{\mathcal{K}(s^{+},q)+\mathcal{K}(s^{-},q)}\right],$
ensuring that our re-ranker effectively distinguishes and prioritizes the most relevant document chunks.
The DRR pipeline is iteratively implemented as follows:
  1. (1)
    Retrieval and Annotation: Extensive results are retrieved using FAISS in the Multi-Path Retrieval pipeline (Section 3.2) for a rich set of candidate documents in each query. Retrieved documents are annotated to create preference samples for DPO, distinguishing between relevant and irrelevant matches. The annotated pairs are then organized into a structured dataset with the balance between positive and negative samples maintained to prevent training bias.
  2. (2)
    Evaluation: Multiple evaluation metrics, including Normalized Discounted Cumulative Gain (NDCG) and Mean Reciprocal Rank (MRR) (Schwartz, 2021), are employed to verify the model’s ability to determine document relevancy and hence improve DRR’s performance.
  3. (3)
    Model Adjustment: When the model fails to perform on a new test set, the retrieved documents are re-annotated to better represent the underlying patterns and relationships via a careful review and adjustment of training data. The model is retrained on top of the previous version using the updated annotations, refining the ranking ability. This is repeated until a satisfactory performance on the new test set is attained.
After document re-ranking, the RAG pipeline in FinSage follows the traditional method to handle user queries, retrieve related data from the vector database, and generate answers accordingly (Figure 2).
4. Experiments
We conduct experiments across three settings to evaluate FinSage: (1) assessing the Multi-path Retrieval pipeline, (2) evaluating the Document Re-ranker, and (3) measuring overall answer quality.
4.1. Datasets and Baseline
To evaluate performance, a manually annotated dataset is typically regarded as the “golden dataset,” with high-quality entries serving as the ground truth (klu.ai, [n. d.]). We utilize two such datasets:
FinanceBench
FinanceBench (Islam et al., 2023) is a financial question-answering dataset created to evaluate the performance of LLMs on open-book tasks. It includes 150 open-source questions relating to traded companies and provides raw financial documents in PDF format containing straightforward answers. We use 149 questions from FinanceBench and construct a single vector store by processing 83 documents that contain correct answers for these questions with FFP. A fixed length of 256 tokens is adopted when separating semantic segments into chunks. We compare our proposed RAG system with three baseline methods, namely Islam et al. (2023), Jimeno-Yepes et al. (2024) and Setty et al. (2024) with their best-matching variants Shared Vector Store, Base256, and Reranker, respectively. These variants generally follow a setup similar to ours, utilizing a single vector store for all filings in FinanceBench, a chunk size of 256 tokens, as well as the retriever before generating answers. The baseline reranker is combined with a cross-encoder reranker which best aligns with our configuration.
In-depth Single Company Dataset
Apart from FinanceBench, we construct a custom QA dataset using 12 filings from a single company. These filings are segmented into a vector database, with each chunk fixed at 200 tokens. We manually compile 75 question-answer pairs from SEC 6-K and F-1 filings, associating each question with an average of 8.7 relevant text chunks containing the most pertinent passages. The questions cover various aspects of Lotus Technology Inc., including products, sales data, history, and corporate strategy. To ensure accuracy, we engage domain experts to review the manually annotated chunks, establishing them as the ground truth. The complexity of these questions often necessitates aggregating information from multiple documents or sections, making it challenging to locate answers within a single chunk. We refer to this dataset henceforth as the Company dataset.
4.2. Experiment Settings
4.2.1. MPR settings
We present the experiment settings for the Multi-Path Retrieval (MPR) pipeline, following the order of its retrieval workflow. After the query process, FinSage selects the top-$K$ results from the vector database to match the query.
In our HyDE retriever, we train a Qwen‑2.5‑7B‑Instruct to follow the HyDE instructions and acquire finance knowledge from the 72B model. To synthesize the training dataset, we first generate question-document pairs using Qwen-2.5-72B-Instruct and select the top 2,000 pairs based on similarity scores. Two hypothetical chunks are generated for each pair, encapsulating essential domain-specific nuances. After training, the model produces three hypothetical answers for each question during testing, which are then used to retrieve the corresponding set of top-$K$ chunks.
For the metadata retriever, the process selects the top-$K$ metadata entries and includes all chunks from the corresponding semantic segments as candidates. Accordingly, to maintain a consistent number of retrieved chunks across different methods, we adopt varying top‑K values.
Finally, unique chunks are collected across different retrieval methods, after which we select $\tau_{\text{exp}}=0.85$ for the chunk expansion threshold with a limit of $\mathcal{K}=5$, to include the neighbor chunk with sufficient similarity.
4.2.2. DRR settings
We assess the impact of Document Re-ranker on retrieval effectiveness using different configurations (Top $K$ + Different Re-ranker), as detailed in Figure 5. Here, Top $K$ denotes the number of top-reranked chunks by the re-ranker. For re-ranking, we compare our fine-tuned DRR, trained on single-company financial filings, against BAAI/bge-reranker-v2-gemma, a general re-ranker. Our model is fine-tuned using Direct Preference Optimization (DPO) with a contrastive learning objective. The training dataset comprises annotated query-document pairs, each containing one relevant chunk and multiple irrelevant ones. This enables the model to adapt to SEC filings’ terminology and structure, ranking relevant chunks higher for improved retrieval precision. The experiment aims to evaluate whether a domain-specific re-ranker enhances retrieval effectiveness.
4.2.3. Response settings
We evaluate the end-to-end question answering (QA) performance of FinSage by comparing it against question answering approach proposed in  (Setty et al., 2024; Jimeno-Yepes et al., 2024; Islam et al., 2023) on the FinanceBench dataset. Additionally, we conduct QA accuracy evaluation on our in-depth company dataset. For FinanceBench, we construct a single vector store (one store for all filings, chunk size = 256) by processing financial PDF documents using the Financial Filings Preprocessing Pipeline (FFP) and apply our multi-path retriever combined with a fine-tuned document re-ranker to retrieve relevant chunks before generating answers. This setup on FinanceBench resembles the setup of Reranker (Setty et al., 2024), Base256 (Jimeno-Yepes et al., 2024), and Shared Vector Store (Islam et al., 2023). In addition, we conduct a separate end-to-end evaluation on the Company dataset, where we measure QA accuracy by comparing generated answers against ground truth answers. For response generation, we set $K=5$ for each sub-query to efficiently manage LLM token consumption. This threshold is justified by empirical evidence indicating that retrieving five chunks is generally sufficient to generate effective answers (Liu et al., 2023).
Table 1. Retrieval Performance on the company dataset. “Avg Num Retrieved” represents the average number of chunks in the retrieved set. Recall, Precision, and F1 scores are calculated based on the relevant chunks retrieved. The bold denotes the best Recall, Precision and F1 produced from FinSage.
Category Method Avg Recall Avg Precision Avg F1 Avg Num Retrieved
FAISS Retrieval FAISS (Baseline) 0.8194 0.0941 0.1646 75.6
 + Bundle Expansion (Exp) 0.8103 0.0999 0.1733 70.25
BM25 Retrieval  + BM25 0.8452 0.1147 0.1949 66.24
Metadata Retrieval  + Metadata 0.8573 0.1198 0.2037 64.87
HyDE Retrieval  HyDE-1: HyDE(Qwen7B) 0.8228 0.1076 0.1830 69.09
 HyDE-2: HyDE(Qwen7B-SFT) 0.8567 0.1072 0.1831 72.92
 HyDE-3: HyDE(Qwen72B) 0.8323 0.1078 0.1844 69.77
FinSage  + BM25+Metadata+HyDE-2 + Exp 0.9251 0.1272 0.2156 68.75
4.3. Evaluation
MPR Evaluation
Our MPR pipeline is evaluated using various combinations of retriever paths by analyzing document chunk recommendation metrics (Recall, Precision, F1) on the in-depth dataset. By introducing additional retriever paths, we expect improvements in recall rates due to a more diverse information coverage (Zhu et al., 2024).
Figure 4. The percentage of relevant chunks selected by retrieval methods relative to the total number of retrieved relevant chunks. The numbers in the overlapping regions indicate the percentage of relevant chunks retrieved by two or more retrieval methods.
[…]
Overlap of Positive Chunks across retrieving methods.
DRR Evaluation
We assess the effectiveness of the Document Re-ranker by comparing the retrieved chunks with the manually annotated ground truth dataset. The re-ranker prioritizes and selects up to $K$ chunks per sub-query, where $K$ is an adjustable parameter. We then measure the re-ranking performance with Precision, Normalized Recall, Mean Reciprocal Rank (MRR), and Binary Normalized Discounted Cumulative Gain (nDCG). To ensure a fair assessment, we introduce an adjusted Normalized Recall metric:
Normalized Recall= correctly retrieved $\text{Normalized Recall}=\frac{\text{ correctly retrieved chunks}}{\min(\text{ground truth chunks},K)}\;.$
We cap denominator to $\min(\text{ground truth chunks},K)$, making it an upper bound for recall calculation. Traditional recall would underestimate and unfairly penalize the model if we were to divide by the actual number of ground truth chunks, despite the reranker functioning optimally within its retrieval limit.
Response Evaluation
We compare between the quality of system-generated responses and expert-provided answers through automatic evaluation using GPT-4o. Specifically, we use the CorrectnessEvaluator module from LlamaIndex, which utilizes GPT-4o to systematically score generated responses based on alignment of FinSage’s outputs with ground truth reference answers. This follows prior work (Moore et al., 2023; Naismith et al., 2023; Zheng et al., 2023) which utilize language models to judge the quality of system-generated responses. To complement automatic evaluation, our team manually labeled the responses after careful discussion. Human evaluation focuses on factual correctness, coherence, and completeness, however, minor factual inconsistencies may be tolerated if they do not alter the overall meaning or intent of the response. Details are provided in the appendix. These two metrics together provide a measurement to assess the FinSage’s ability to generate responses that align with the annotated ground truth answers both textually and semantically.
Figure 5. Performance of different re-rankers under different retrieval settings.
5. Results and Analysis
5.1. MPR Results
Table 1 demonstrates FinSage’s superior performance through its novel multi-path retrieval architecture. While traditional single-path approaches like FAISS and BM25 show decent performance in isolation, they each capture only limited aspects of the complex financial domain based on superficial semantic matching. FinSage’s innovative combination of multiple specialized retrievers, particularly its integration of metadata-aware retrieval with fine-tuned HyDE hypothetical document generation, allows it to comprehensively capture both the structural and semantic aspects of financial documents. This synergistic approach enables FinSage to maintain high precision while significantly improving recall.
Our experimental results reveal the fine-tuned Qwen2.5-7B-SFT model surpasses a significantly larger Qwen2.5-72B counterpart across key metrics, particularly in recall performance. This highlights the critical role of fine-tuning with high-quality domain-specific data, reinforcing the necessity of FinSage for precise and accurate question-answering in the complex structure of financial filings. The strong performance of our smaller fine-tuned model aligns with emerging research demonstrating that well-adapted smaller language models can match or exceed the performance of larger models in specialized domains (Bucher and Martini, 2024), while offering substantial computational efficiency advantages.
To better understand the uniqueness of each retrieval method, we analyze their overlap on the Company dataset. As shown in Figure 4, while each method uniquely contributes to retrieving positive chunks, there is significant overlap, suggesting that hybrid approaches could leverage their combined strengths. Notably, FAISS and HyDE retrieve highly similar positive chunks, with an overlap rate of $52.7\%$ $(15.3\%+12.9\%+15.3\%+9.2\%)$, compared to their individual recovery rates of $62.9\%$ and $69.6\%$. This overlap likely arises because HyDE generates hypothetical documents by referencing the original document, preserving much of its context.
5.2. DRR Results
We compare re-rankers’ performance in Figure 5, where we compare our trained re-ranker with bge-reranker-v2-Gemma. These results demonstrate that our Document Re-ranker significantly and consistently outperforms the general re-ranker across all settings and metrics in top-5 and top-10 configurations. Notably, our trained re-ranker retrieves more relevant chunks, evidenced by an approximate 15% improvement in recall across all setups. Moreover, performance gains are not limited to retrieving relevant chunks, it also effectively filters out irrelevant information, leading to a substantial precision boost from 5.6% to 18.5%.
It is worth noting that precision, MRR, and binary nDCG all decline as more chunks are extracted by the reranker: precision drops from 78.8% to 38.8%, MRR from 75.4% to 20.7%, and binary nDCG from 85.3% to 53.5%. This consistent trend across benchmark scenarios indicates that extracting additional chunks increases the burden on the re-ranker, resulting in less optimal ranking in chunk selection during re-ranking. Moreover, the rapid degradation in these metrics between Top-10 and Top-15 aligns with previous studies suggesting that retrieving only 5 to 10 chunks is generally sufficient for producing effective answers (Liu et al., 2023).
5.3. Response Results
Table 2. Performance of End-To-End QA. The bold is from our implementation, which has the top LLM and Manual scores. Results with * are from our experiments, while others are from the original papers.
Dataset Method LLM Manual
FinanceBench Islam et al. (2023) - 0.1900
FinanceBench Jimeno-Yepes et al. (2024) 0.3262 0.3688
FinanceBench Setty et al. (2024) 0.2560 -
FinanceBench FinSage 0.4966 0.5705
Company FinSage 0.8533 0.8800
Table 2 shows performance on the FinanceBench and Company datasets. FinSage achieves a significant improvement over prior methods on FinanceBench, with an LLM accuracy of 49.66%, outperforming Jimeno-Yepes et al. (2024) (32.62%) and Setty et al. (2024) (25.60%, aligns most closely with our configuration). Similarly, manual evaluation also reflects an accuracy of 57.05%, substantially higher than the closest competitor  Jimeno-Yepes et al. (2024) (36.88%). These results indicate that FinSage’s retrieval and ranking mechanisms contribute significantly to improved answer quality. In the Company dataset, FinSage also demonstrates strong performance, which achieves an LLM accuracy of 85.33% and a manual accuracy of 88.00%. The consistently higher scores across both datasets indicate that FinSage is particularly adept at processing structured financial filings with clear contextual information. By retrieving semantically relevant passages across multiple documents, FinSage significantly improves response accuracy, making it highly effective for complex financial question-answering tasks.
6. Conclusion
In this paper, we propose FinSage, a RAG solution for the question-answering task on financial filings. FinSage comprises a multi-modal data pre-processing pipeline, a domain-aware multi-path retrieval system specifically designed for financial data, and a domain-specialized re-ranking module fine-tuned using DPO. Our system demonstrated state-of-the-art performance on expert-curated questions derived from financial filings. Moreover, FinSage has been deployed as a financial question-answering agent in online meetings organized by affiliated entrepreneurs.
Appendix A Data Preprocessing Details
To effectively convert PDF files into structured data, we utilize the Mineru tool  [^1] https://github.com/opendatalab/MinerU, which can extract content from PDFs and export it in various formats. The preprocessing workflow is described below.
A.1. PDF to JSON Conversion Process
We can utilize the Mineru to to process PDF files with the following command:
magic-pdf -p {some_pdf} -o {some_output_dir} -m auto
Users may need to specify the input PDF file and the output directory. The tool generates the following output files:
  • some_pdf.md: Markdown representation of the PDF content.
  • images/: A directory containing extracted images, including tables.
  • some_pdf_layout.pdf: A visualization of the document’s layout.
  • some_pdf_middle.json: Intermediate processing results in JSON format.
  • some_pdf_model.json: Model inference results in JSON format.
  • some_pdf_origin.pdf: The original PDF file.
  • some_pdf_spans.pdf: A PDF file showing the bounding box positions at the smallest granularity.
  • some_pdf_content_list.json: A structured JSON file representing the document content in reading order.
Among these files, some_pdf_content_list.json stores the document content into blocks, each of which contains the following fields: type, text, text_level, and page_idx. The type field differentiates different content types, such as:
- text: Regular text blocks. - table: Table elements. - img_path: Image paths referencing extracted tables.
A.2. Examples of JSON Structure
Text Block Example:
[…]
Table Block Example:
[…]
Image Block Example:
[…]
A.3. Image Processing and Text Parsing
To extract table information, the Mineru2Base tool leverages the GPT-4o API to convert table images into text. The system iterates through each document block in some_pdf_content_list.json . For blocks containing an img_path , the tool retrieves the preceding and succeeding blocks as context. The table image is then converted to Base64 format and passed to the GPT-4o API, which generates a structured textual representation.
A.4. Embedding Model
Two embedding models are used in this project, each serving a distinct purpose: TF-IDF for text deduplication and BGE-M3 for storing text in the database.
For the deduplication task, we employ the TF-IDF tool to convert the text into vector representations. This approach helps to identify and remove redundant text segments by comparing their vector similarities. The model uses term frequency and inverse document frequency to generate vectors, which are then compared using cosine similarity to detect high similarity between text segments. Text with similarities above a predefined threshold is removed, reducing redundancy.
For storing text in the Chroma database, we use the BGE-M3 embedding model. This model, trained on large-scale datasets, transforms text into high-dimensional vectors that capture the semantic meaning of the content. These vectors are then stored in the database, enabling efficient indexing and retrieval for future similarity searches and semantic analysis.
A.5. Duplicate Removal
To remove redundant content, a de-duplication step is performed using text similarity measurements. The method includes:
1.Chinese word segmentation using the Jieba library. 2.ransforming text into a term-frequency matrix via PyTorch, where the text is vectorized using a TF-IDF approach implemented with torch for efficient matrix computation. 3.Computing cosine similarity between document blocks using PyTorch’s cosine similarity function.
If the similarity score exceeds 0.7 (default threshold), the duplicate block is removed.
A.6. Coreference Resolution
Coreference resolution is an important task in natural language processing, aiming to identify the entities that pronouns and other referring expressions point to within a text. In simple terms, it analyzes pronouns (such as “he,” “it,” “we,” etc.) and determines the specific object or entity they refer to, making the text more coherent and clear. The relationship between pronouns and their referred entities can significantly affect comprehension, and thus coreference resolution plays a crucial role in improving text readability and reducing ambiguity.
In practical implementation, the algorithm traverses a JSON file containing multiple chunks, each with a title and content field. For each chunk, the coreference resolution follows a specific process: it first searches up to four preceding chunks as references. The search is limited to preceding chunks that share the same title as the current one. If the previous chunk belongs to a different title, it is considered to have low relevance to the current chunk and is thus not considered as a reference.
In this way, the coreference resolution algorithm effectively scans the related content within the text, using the GPT-4o-mini API to resolve pronoun references and replace them with specific entities. This approach not only improves the clarity and coherence of the text but also reduces the ambiguity that pronouns may cause, ensuring the accurate transmission of information.
Detail prompts provided to GPT-4o-mini during the anaphora resolution step are shown in Figure 6:
Figure 6. The instruction prompts for anaphora resolution.
A.7. Title Summarization
Each block receives a summary stored in the title_summary field. The process involves:
1. Generating a concise summary for each block. 2. Aggregating summaries to form a higher-level title summary. 3. Storing the final summary in title_summary.
For example, the original JSON format:
After summarization:
A.8. Text Segmentation
To maintain manageable block sizes, text segmentation is applied. The segmentation rules are:
- If a block has fewer than 200 characters, subsequent content is appended until it exceeds 200. - If a block exceeds 200 characters, the last sentence is moved to the next block. - Segmentation is restricted within the same title to prevent cross-title segmentation.
Through this structured preprocessing pipeline, PDF content is efficiently converted into JSON format, ensuring accurate extraction of text, tables, and other elements. This process facilitates subsequent analysis, model applications, and further information retrieval.
A.9. Loading Chunk into a Vector Database
In this study, text data is loaded into a Chroma-based vector database, using the BGE-M3 embedding model to vectorize the text. The process begins by reading a series of JSON-format files from a specified directory, each containing metadata related to page content and the text itself. Metadata such as page range and publication date are extracted and processed along with the corresponding content. To avoid duplicate data, each piece of text is uniquely identified by calculating its SHA-256 hash, and in case of duplicate content, the system updates the record with the latest version based on the publication date.
Next, the text and associated metadata are stored into the Chroma database in batches. During the storage process, the system also maintains the sequential relationship between documents by adding “previous” and “next” chunk identifiers for texts within the same file. Finally, to further optimize information retrieval, a BM25 index is built based on the text data stored in the Chroma database, improving retrieval efficiency in subsequent searches.
Through this process, we can effectively convert large amounts of text data into vectors and provide efficient database support for future retrieval and analysis.
Appendix B Reranker Finetune Details
B.1. Data Preparation
The training data is stored in a JSONL format, where each entry contains the following fields: query, pos, neg, and optionally, pos_scores, neg_scores, and prompt. The key fields are as follows:
  • query: The query string.
  • pos: A list containing one positive candidate related to the query.
  • neg: A list containing multiple negative candidates unrelated to the query.
  • pos_scores (optional): A list of scores corresponding to the positive candidates.
  • neg_scores (optional): A list of scores corresponding to the negative candidates.
  • prompt (optional): A prompt string used for additional customization of the training data.
The minimal required data format is as follows:
In each labeled sample, the pos field is structured as a list containing only a single positive instance. This positive instance typically represents the most relevant or ideal answer in relation to the query, thereby providing the model with a clear and representative positive reference. In contrast, the neg field consists of a list of multiple negative candidate strings, each reflecting content that is either irrelevant or only weakly related to the query. This data construction facilitates a significant enhancement of the model’s sensitivity and discriminative capability in assessing relevance, achieved through contrastive learning between positive and negative samples during the fine-tuning of the reranker model.
Example of the labeled training data:
The above format is used for fine-tuning a reranker model, where the goal is to reorder the candidate passages based on their relevance to the query.
B.2. Selection of Positive and Negative Samples
In this study, the selection of positive and negative samples is governed by a systematic and rigorous procedure, meticulously designed to ensure that the chosen samples effectively facilitate both the training and subsequent optimization of the model. The process comprises several carefully orchestrated stages, which are detailed as follows:
Initially, a comprehensive set of questions was curated to represent a wide array of potential query scenarios that the model might encounter in real-world applications. These questions were selected to capture the inherent diversity and complexity of user inquiries, thereby ensuring that the dataset is both representative and robust. For each question, a dual retrieval strategy was implemented using the Faiss and BM25 algorithms within the Chroma vector database. Faiss, renowned for its efficiency in high-dimensional vector similarity search, and BM25, a well-established probabilistic retrieval model, complement each other by capturing both semantic nuances and lexical patterns. This combined approach ensures that each query retrieves a diverse set of candidate answers—typically exceeding twenty per question—thus providing a rich pool of potential responses with varying degrees of relevance and quality.
Subsequent to the retrieval phase, the candidate results underwent an exhaustive manual labeling process. Expert annotators meticulously evaluated each candidate to assess its capacity to accurately and comprehensively address the corresponding question. Candidates that met stringent criteria for precision, relevance, and completeness were designated as positive samples. These positive samples are intended to exemplify the optimal answers, providing a clear and reliable reference for the model during training. Conversely, any candidate that failed to deliver an effective or accurate answer was classified as a negative sample. This deliberate bifurcation into positive and negative samples ensures that the model is exposed to a distinct contrast between high-quality and substandard responses, thereby enhancing its ability to discriminate between relevant and irrelevant content.
To ensure fairness, consistency, and reproducibility throughout the sample selection process, each question was processed using an identical evaluation protocol. Strict adherence to uniform criteria and standardized methods during the manual labeling phase served to minimize potential biases that might otherwise arise from automated selection procedures. This rigorous approach not only guarantees the high quality of both positive and negative samples but also underpins the overall robustness of the training data. Ultimately, this meticulous curation process is instrumental in bolstering the model’s performance, particularly in its capacity to accurately discern and rank relevant content in response to diverse queries.
B.3. Training Parameters
During the fine-tuning of the reranker model, Low-Rank Adaptation (LoRA) was used to reduce the number of parameters that need to be updated during training, thus enhancing training efficiency. Specifically, LoRA was enabled, with the rank set to 32 and the scaling factor set to 64. The learning rate was set to 1e-4 to ensure stable training and prevent gradient explosion. To save memory and improve stability, the batch size per device was set to 1. Additionally, gradient accumulation was enabled, with gradients updated after processing two batches, allowing for more efficient training while maintaining a small batch size. The model was trained using 10 epochs, and mixed precision training was employed with the bf16 setting to further optimize performance and memory usage.
Appendix C RAG Evaluation
To evaluate the quality of the Retrieval-Augmented Generation (RAG) model, this study employs two assessment methods: large model evaluation and human evaluation. First, a representative set of questions is collected, covering a range of query scenarios that the model may encounter. For each question, the entire RAG process is executed, including retrieval, generation, and ranking steps, ultimately producing the model’s answers. To facilitate subsequent analysis and evaluation, these generated results are stored in a standardized JSON file format, ensuring traceability and structured data. Through this approach, we can comprehensively assess the performance of the RAG model, evaluating its performance under automated evaluation metrics while also incorporating human review to further analyze the accuracy and relevance of its generated content.
Figure 7. Predefined evaluation prompt template used in LlamaIndex’s CorrectnessEvaluator.
C.1. Environment Setting
Our experiments were conducted on a NVIDIA A100-SXM4-80GB server, where we deploy the Qwen2.5-72B-Instruct-AWQ 4bit model. The model was configured using the following command:
Table 4. Model deployment parameters using VLLM.
Parameter value
max-model-len 6144
gpu-memoty-utilization 0.7
engorce-eager True
swap-space 36
C.2. LLM Evaluation
In this approach, we utilize the open-source tool llamaindex  (Liu, 2022) to systematically evaluate RAG model outputs by loading two JSON files: one containing generated QA pairs from the RAG model, and the other providing reference answers with contextual data in content_list. Each question is mapped to its reference information, and a Response object is constructed by combining the RAG-generated answer with its concatenated context. We then employ the CorrectnessEvaluator module (backed by a GPT-4o api) to assess each Response, returning a binary passing flag along with detailed feedback.
To facilitate the evaluation, we integrate the correctness assessment module from LlamaIndex’s CorrectnessEvaluator. This module assigns a numerical score (ranging from 1 to 5) to each response based on alignment with reference answers, using a predefined evaluation template. A threshold-based scoring mechanism determines whether a response meets the required correctness level. The evaluator asynchronously processes input queries, system responses, and reference answers using a structured prompt, ensuring consistency and scalability in judgment. The predefined evaluation prompt template encapsulated in the module is as follows:
During the evaluation, logs and statistics are recorded, including total queries processed, successful evaluations, and skipped items, while exceptions are noted and the evaluation continues for remaining data. Finally, the results—comprising the question, generated answer, reference answer, evaluation status, and feedback—are stored in a structured JSON file for further analysis, providing a robust, LLM-driven metric for measuring and tracing the fidelity of RAG-generated responses.
An example of the result is as below:
Figure 8. Example question answering results
C.3. Re-ranker Evaluation
Binary Normalized Discounted Cumulative Gain (BnDCG) measures ranking quality by applying position-aware discounting with logarithmic scaling. Since our manually annotated dataset lacks explicit ranking scores for ground truth chunks, we adopt a binary relevance model (1 for relevant, 0 for irrelevant). DCG and its ideal counterpart (IDCG) are computed as:
(3) $\text{DCG}=\sum_{i=1}^{n}\frac{rel_{i}}{\log_{2}(i+1)}$
where $rel_{i}$ is the binary relevance score of the retrieved chunk at position $i$, taking values of either 1 (relevant) or 0 (irrelevant).
The Ideal DCG (IDCG) is computed by sorting all relevant chunks to the top positions:
(4) $\text{IDCG}=\sum_{i=1}^{n^{*}}\frac{1}{\log_{2}(i+1)}$
where $n^{*}$ is the total number of relevant chunks in the dataset, and $rel_{i}=1$ for all positions $i$ in the ideal ranking, ensuring that all relevant chunks appear first. Finally, BnDCG is given by:
(5) $\text{BnDCG}=\frac{\text{DCG}}{\text{IDCG}}$
ensuring a normalized ranking score between 0 and 1, where 1 represents a perfect ranking.
Table 5. Performance comparison of different re-rankers under different retrieval settings.
Method Precision Normalized Recall MRR Binary nDCG
5 Candidate Results per Retriever (R=5):
Top-5 BGE 0.5913 0.6043 0.6570 0.7540
Top-5 trained Reranker 0.7324 0.7456 0.7078 0.8124
Top-10 BGE 0.3771 0.6881 0.3854 0.6011
Top-10 trained Reranker 0.4424 0.7783 0.3338 0.6097
10 Candidate Results per Retriever (R=10):
Top-5 BGE 0.6028 0.6130 0.6540 0.7638
Top-5 trained Reranker 0.7878 0.7910 0.7545 0.8533
Top-10 BGE 0.4133 0.5985 0.4533 0.6285
Top-10 trained Reranker 0.5657 0.8196 0.5155 0.6958
Top-15 BGE 0.3318 0.6889 0.3974 0.5680
Top-15 trained Reranker 0.3883 0.8064 0.2071 0.5354
Table 6. Performance of End-To-End Answer (ROUGE scores included)
Dataset/Method LLM Mannual ROUGE1 ROUGE2 ROUGEL
FinanceBench/FinSage 0.4966 0.5705 0.3319 0.1542 0.3002
Company/FinSage 0.8533 0.8800 0.7213 0.5576 0.6657
Figure 9. Performance comparison of different re-rankers under different retrieval settings when limiting the retrieval bundle to 5.
Appendix D Additional Results
D.1. Re-ranker Results
We also assess the impact of Document Re-ranker on retrieval effectiveness using different $R$, as detailed in Figure 5. $R$ denotes the number of retrieved chunks by the retrievers. The retrieval system under this assessment integrates four retrievers—FAISS (baseline), BM25, Metadata, and HyDE—retrieving at least $4\times R$ chunks. While increasing $R$ enhances the chances of retrieving relevant chunks, it also intensifies the re-ranker’s challenge of filtering out irrelevant ones.
D.2. Final Response
As part of our evaluation, we also compute standard metrics as shown in Table 6, including ROUGE1, ROUGE2 and ROUGE-L scores, to assess the alignment between system-generated responses and reference answers. However, these scores often fall short in capturing the semantic equivalence of different responses, which leads to potentially misleading evaluations. Given such limitation, we rely primarily on LLM-based evaluation using GPT-4o and human labeling,